# GraspMolmo: Generalizable Task-Oriented Grasping via Large-Scale Synthetic Data Generation

### v2

@@ Abstract
We present GraspMolmo, a generalizable open-vocabulary task-oriented grasping (TOG) model. GraspMolmo predicts semantically appropriate, stable grasps conditioned on a natural language instruction and a single RGB-D frame. For instance, given “pour me some tea,” GraspMolmo selects a grasp on a teapot handle rather than its body. Unlike prior TOG methods, which are limited by small datasets, simplistic language, and uncluttered scenes, GraspMolmo learns from PRISM, a novel large-scale synthetic dataset of 379k samples featuring cluttered environments and diverse, realistic task descriptions. We fine-tune the Molmo visual-language model on this data, enabling GraspMolmo to generalize to novel open-vocabulary instructions and objects. In challenging real-world evaluations, GraspMolmo achieves state-of-the-art results, with a 70% prediction success on complex tasks, compared to the 35% achieved by the next best alternative. GraspMolmo also successfully demonstrates the ability to predict semantically correct bimanual grasps zero-shot. We release our synthetic dataset, code, model, and benchmarks to accelerate research in task-semantic robotic manipulation, which, along with videos, are available at this URL.
@@
Figure 1: A. We introduce a large dataset of task-semantic-annotated tasks in virtual scenes with a large variety of target objects, which we name Purpose-driven Robotic Interaction in Scene Manipulation. B. We use PRISM to train GraspMolmo, which represents grasps as points and matches them to grasp proposals. C. We show strong transfer and generalization in real-world complex scenes with interesting task semantics using a 7-DoF Franka FR2 arm, and additionally show zero-shot adaptation to semantically-appropriate bimanual grasping.
@@ 1 Introduction
Robotic grasping has come a long way, with models now capable of robustly predicting stable grasps for a wide range of objects. Yet most existing methods operate in an object-centric fashion: they predict grasps that are stable for a given object but oblivious to the task at hand. In the real-world, however, how an object should be grasped depends crucially on what the robot is trying to do with it.
@@
Recent work has introduced datasets and benchmarks [2, 3, 4], but they are limited in realism and diversity. Some contain relatively simple scenes without clutter and lack diverse object categories [2]. Others often templatize (e.g., “grasp the mug to pour” or “grasp the pan to clean”). These simplifications fail to capture the complexity of real-world environments and the wide variety of language humans use to specify tasks, such as “mince some garlic” or “do the dishes”. Furthermore, many models require pre-segmented point clouds [4] or multi-view observations [3], restricting real-time deployment.
To address these limitations, we propose GraspMolmo: a generalizable task-conditioned grasping model trained entirely on synthetic data. To overcome the brittleness of prior datasets, we create Purpose-driven Robotic Interaction in Scene Manipulation (or PRISM for short), a large-scale suite of task-grasp pairs with 10,000 scenes, 2,356 object instances, and task descriptions ranging from simple (“cut the apple”) to compositional or nuanced (“mince some garlic for a salad”). This dataset spans diverse cluttered scenes with realistic textures, occlusions, and multiple objects per task.
Using this dataset, we fine-tune Molmo [5], a recent open-weight vision-language model (VLM), to predict 6-DoF grasps from a single RGB-D frame and a task instruction. By grounding task semantics in the visual scene, GraspMolmo learns to disambiguate grasps that would otherwise be indistinguishable from a purely geometric perspective. GraspMolmo achieves state-of-the-art results on the TaskGrasp benchmark (76.7% versus the next best prior work at 72.3%). GraspMolmo also achieves significant performance boost on our benchmark of simulated cluttered scenes (62.5% versus 40.0%). Naturally, simulated evaluations are only a proxy, but we further demonstrate GraspMolmo transfers zero-shot to the real world in cluttered scenes, including objects and tasks completely unseen in training (70.4% prediction and 61.1% overall success rates). GraspMolmo also zero-shot transfers to semantically-appropriate bimanual grasping in qualitative testing. We will release our dataset, model, code, and benchmarks to enable further research in task-semantic manipulation.
@@ 2 Related work
Task-oriented grasping. The field of Task-Oriented Grasping (TOG) addresses how robots should grasp objects based on intended tasks rather than just stability. Recent approaches have shown diverse strategies with specific limitations. GraspGPT [6] and FoundationGrasp [7] are most similar to our method, but due to their reliance on TaskGrasp, have limited pretraining with simple scenes and tasks. LERF-TOGO [8] uses language-embedded radiance fields to localize grasp points in a scene, but has costly requirements such as multiple views and long inference times, which make it impractical for use. RTAGrasp [9] transfers grasping behavior from human demonstrations but depends heavily on task-grounded data. Methods such as CROG and GraspCLIP [10, 11] represent a family of methods that perform inference directly on the image, outputting per-pixel predictions for grasp poses. These methods are limited to predicting only 4-DoF grasps, and strictly require a top-down camera view. TOGNet [12] extracts “patches of interest” in the image for grasping, but only accepts object-level rather than task-level grounding, resulting in coarse semantics that are ill-suited for diverse tasks. Underpinning many of these works, the TaskGrasp benchmark [13] provides evaluation standards and training data, but features simple scenes with limited task diversity. Our work addresses these collective limitations through comprehensive synthetic data generation with complex scenes, diverse and natural task specifications, and large data scale.
@@
Vision-language models in robotics Recent years have witnessed significant progress in leveraging the strong vision-language commonsense capabilities of foundation multimodal language models (MLMs)  [14, 15, 16, 17] for taking actions in the physical world [18, 19, 20, 21, 22]. Some studies employ pre-trained MLMs directly in a prompt-based manner [23, 24, 25, 26], while others further fine-tune or instruction-tune models for specific capabilities, including action policies [27, 19, 21, 28], physics [29], navigation [30, 31], and spatial reasoning [32, 33, 34]. Our approach aligns with the latter, fine-tuning a state-of-the-art MLM [15] on task descriptions and grounding points specifically for object grasping tasks. We advance beyond concurrent methods that separately address physics or multi-step reasoning limited to judging object relationships [33, 29] or precise pointing [32] without multi-step reasoning, by integrating both aspects (e.g., needing to reason about the precise point to grasp a pot to pour tea). In contrast to vision-language-action models that directly predict low-level controller actions [22, 21, 35], our method predicts grasp points in image space, enhancing generalizability and avoiding embodiment-specific constraints, thereby serving as a robust foundation model deployable across diverse hardware. While works like [31, 36, 32, 30] also predict points or high-level plans that are generalizable, they mainly focus on navigation and pick-and-place, while we focus on precise object manipulation, which is more difficult [36]. We also take inspiration from concurrent works [33, 37, 31] that show simulations as a scalable data source for real-world generalization. However, we generate large-scale, task-oriented grasping data that focuses on complex task-dependent object manipulation that goes beyond simpler QA or navigation tasks in these works.
@@
Object affordances in robotics. Affordances describe the functional properties of objects—how they can be manipulated—linking perception to action beyond appearance. Affordance prediction has driven advances in learning-based 6-DoF grasping [38, 13, 39] and stable object placement [40, 41, 42, 32]. Researchers have explored various representations, including part-level segmentation [43], dense visual descriptors [44], and keypoint-based encodings [45, 46, 47]. At the policy level, affordances are often learned implicitly from demonstrations and generalized using rich visual features, as seen in TransporterNet [40], CLIPort [48], and recent works leveraging vision-language models (VLMs). However, many of these approaches rely on pretrained VLMs for visual grounding without incorporating deeper semantic reasoning, often resulting in less accurate or unstable point detection. In contrast, our approach instruction-tunes Molmo, a VLM enhanced for semantic understanding, to generate task-aware, stable grasp points through contextualized spatial reasoning.
@@ 3.1 Generation of PRISM-Train and PRISM-Test
Our dataset employs assets from ShapeNet-Sem [49, 51] with grasps derived from ACRONYM [50]. We selected 91 object classes for use in data generation, enumerated in the supplement. Scene composition was performed using SceneSynthesizer [52], which allows for procedural generation of realistic object arrangements. We implemented additional lighting variations and camera randomization to increase the visual diversity of the training examples. As illustrated in Figure 2, we generated a comprehensive collection of diverse scenes spanning different object categories, lighting conditions, and spatial arrangements. Each scene was rendered from 10 distinct camera viewpoints to ensure robust training data that captures objects from multiple perspectives. This multi-view approach is particularly important, as grasp affordances often depend on the viewpoint, which can vary wildly in the real world. In our processing pipeline, we carefully track grasp visibility and only include grasps that are clearly visible in the specific renders for later stages of the learning process, ensuring that our model learns only from valid examples.
@@
Generating diverse task-grasp annotations at large scale is intractable, largely due to combinatorics, as every task-object-grasp triple must be annotated. In other words, annotating $G$ grasps per object for $O$ objects across $T$ tasks is $\mathcal{O}(TOG)$. Our insight is that we can drastically reduce the effort required to generate task-grasp annotations by exploiting common structure. Namely, we note that many different tasks might require grasping an object in the same way. Concretely, pouring from a mug and drinking from it both require grasping the handle in the same manner. Therefore, we use grasp descriptions as the bridge between tasks and grasps. Instead of individually annotating each task-object-grasp triple with a binary label, we annotate each object-grasp pair with a natural language description of how the grasp is gripping the object. Separately, for each task we generate descriptions of how an object should be grasped to complete that task. Then, matching tasks to object-grasp pairs reduces to simple text matching. The use of common-sense reasoning and natural language make this task well-suited for automation via LLMs, further aiding in scalability. The use of grasp descriptions decouples tasks from object-grasps pairs, meaning generating annotations is now $\mathcal{O}(T+OG)$, which is significantly more tractable to perform at large scale.
@@
For generating high-quality grasp descriptions, we employed a two-stage approach combining LLMs with human verification. Initially, we used GPT-4o [53] to generate synthetic grasp descriptions. To provide comprehensive visual context for the language model, we created collages containing four different viewpoints of each grasp scenario, enabling the model to consider the full 3D spatial relationship between the gripper and target object, with a sample shown in Fig. 2. To ensure description accuracy and quality, we then engaged human workers through the Prolific platform to verify each description, and provide corrections if necessary. We found that 45% of the synthetically generated grasp descriptions were judged to be accurate while the other 55% required correction, illustrating the importance of human verification for high quality data. Workers were instructed to identify and correct any inaccuracies in the AI-generated descriptions, particularly focusing on spatial relationships, contact points, and approach vectors. This human-in-the-loop verification process significantly improved the precision and natural language quality of our grasp descriptions, creating a dataset that balances the scalability of AI generation with the accuracy of human judgment.
@@
The process to generate semantic grasping tasks, leveraging the available knowledge in GPT-4.1 for a varied and rich set of instructions, involves two steps. Given an object class, we first prompt the LLM to generate two grasp descriptions that are maximally different while remaining plausible for manipulation tasks and avoiding hallucinating features which may or may not exist for an object of the given class. Then, we ask the LLM to generate four valid semantic tasks per grasp while minimizing compatibility of the instruction with the alternative grasp. In total, we generate 728 unique instructions corresponding to two grasps for each of the 91 object classes. For the complete prompts used to generate grasps and task instructions, please refer to the supplement.
@@
Finally, we can match the generated grasp descriptions for each object-grasp pair with that of the generated tasks, which we accomplish using GPT-4o. Concretely, for each object in a scene we first determine the visible grasps, and corresponding grasp descriptions. For each generated task for that object, we then take the corresponding proposed grasp description and ask GPT-4o to determine which annotated object-grasp pair, if any, describes the same grasp. If the LLM determines that one of the existing grasps in the scene is described by the generated description, the task is paired with the object-grasp to create an task-object-grasp triple, and added to the dataset.
@@
To connect grasp data with visual perception, it becomes necessary to map ACRONYM grasps onto the image frame. To do this, we raycast from the grasp center onto the object being grasped, and project the resulting point onto the image plane. This process determines the pixel locations corresponding to each 3D grasp. PRISM-Train  contains 9424 object-centric grasp instances across 2356 unique objects, which get scaled combinatorially via procedural generation. Although we do not use this information for training GraspMolmo , the dataset also includes calibrated 3D grasp poses for all objects, providing ground truth spatial information that can be used for both 2D and 3D perception tasks. The final dataset is 378,844 samples, each of which has a scene render, a task in natural language, a ground-truth semantically appropriate and stable grasp in calibrated camera coordinates, an object-centric spatial description of that grasp in natural language, and a pixel location corresponding to the proper grasp for that task. We present sample generated scenes from PRISM in Figure 4(b), showcasing the diversity in objects, arrangements, and appearances.
@@ 3.2 TaskGrasp-Image
TaskGrasp [2] is a standard dataset in the field of task-oriented grasping, used for training and evaluation. The TaskGrasp dataset consists of partial object point clouds, each annotated with stable grasp poses. Each grasp on each object is annotated with binary labels, indicating whether each grasp satisfies a given task verb (e.g., scoop, cut, stir). However, the use of point clouds makes using image-based models difficult, especially due to limitations in the data such as noisy point cloud reconstructions caused by self-occlusions and segmentation artifacts. Additionally, the use of single verbs and nouns to define a task results in simplistic task conditioning, which insufficiently captures the richness of real-world human instructions.
To address these challenges, we construct TaskGrasp-Image, a new image-based dataset derived from the existing TaskGrasp dataset. Since TaskGrasp fuses multiple RGB-D views to create a single point cloud, we can perform inference on the images themselves instead of the point cloud, which is more suitable for image-based models, and is more realistic due to the absence of fusion, filtering, or segmentation artifacts. As a result, TaskGrasp-Image preserves the ground-truth grasp annotations while placing them in the context of real, unprocessed RGB-D imagery. We defer complete details on deriving TaskGrasp-Image from TaskGrasp to Section A.6.

### v3

@@ Abstract
We present GraspMolmo, a generalizable open-vocabulary task-oriented grasping (TOG) model. GraspMolmo predicts semantically appropriate, stable grasps conditioned on a natural language instruction and a single RGB-D frame. For instance, given “pour me some tea,” GraspMolmo selects a grasp on a teapot handle rather than its body. Unlike prior TOG methods, which are limited by small datasets, simplistic language, and unrealistically simple scenes, GraspMolmo learns from PRISM, a novel large-scale synthetic dataset of 379k samples featuring complex environments and diverse, realistic task descriptions. We fine-tune the Molmo vision-language model on this data, enabling GraspMolmo to generalize to novel open-vocabulary instructions and objects. In challenging real-world evaluations, GraspMolmo achieves state-of-the-art results, with a 70% prediction success on complex tasks, compared to the 35% achieved by the next best alternative. GraspMolmo also successfully demonstrates the ability to predict semantically correct bimanual grasps zero-shot. We release our synthetic dataset, code, model, and benchmarks to accelerate research in task-semantic robotic manipulation, which, along with videos, are available at this URL.
@@
Figure 1: A. We introduce a large dataset of task-semantic-annotated tasks in virtual scenes with a large variety of target objects, which we name Purpose-driven Robotic Interaction in Scene Manipulation. B. We use PRISM to train GraspMolmo, which represents grasps as points and matches them to grasp proposals. C. We show strong transfer and generalization in real-world complex scenes with interesting task semantics using a 7-DoF Franka FR3 arm, and additionally show zero-shot adaptation to semantically-appropriate bimanual grasping.
@@ 1 Introduction
Robotic grasping has progressed significantly, with models now capable of robustly predicting stable grasps for a wide range of objects. Yet most existing methods operate in an object-centric fashion: they predict grasps that are stable for a given object but oblivious to the task at hand. In the real world, however, how an object should be grasped depends on what the robot is trying to do with it.
@@
Recent work has introduced relevant datasets and benchmarks [2, 3, 4], but they are limited in realism and diversity. Some contain relatively simple scenes without any distractors and lack diverse object categories [2]. Others often templatize (e.g., “grasp the mug to pour” or “grasp the pan to clean”). These simplifications fail to capture the complexity of real-world environments and the wide variety of language humans use to specify tasks, such as “mince some garlic” or “do the dishes”. Furthermore, many models require pre-segmented point clouds [4] or multi-view observations [3], restricting real-time deployment.
To address these limitations, we propose GraspMolmo: a generalizable task-conditioned grasping model trained entirely on synthetic data. To overcome the brittleness of prior datasets, we create Purpose-driven Robotic Interaction in Scene Manipulation (PRISM), a large-scale suite of task-grasp pairs with 10,000 scenes, 2,356 object instances, and task descriptions ranging from simple (“cut the apple”) to compositional or nuanced (“mince some garlic for a salad”). This dataset spans diverse, busy scenes with realistic textures and occlusions, and multiple objects per task.
Using this dataset, we fine-tune Molmo [5], a recent open-weight vision-language model (VLM), to create GraspMolmo, enabling it to predict 6-DoF grasps from a single RGB-D frame and a task instruction. GraspMolmo achieves state-of-the-art results on the TaskGrasp benchmark (76.7% versus the next best prior work at 72.3%). GraspMolmo also achieves significant performance boost on our benchmark of simulated realistic scenes (62.5% versus 40.0%). We further demonstrate GraspMolmo transfers zero-shot to the real world in representative household scenes, including objects and tasks completely unseen in training (70.4% prediction and 61.1% overall success rates). GraspMolmo also zero-shot transfers to semantically-appropriate bimanual grasping in qualitative testing. We will release our dataset, model, code, and benchmarks to enable further research in task-semantic manipulation.
@@ 2 Related work
Task-oriented grasping. The field of Task-Oriented Grasping (TOG) addresses how robots should grasp objects based on intended tasks rather than just stability. The TaskGrasp dataset [2] provides training data and evaluation benchmarks for much of the existing TOG literature, but features only simple scenes with limited task diversity and simplistic specification. Leveraging TaskGrasp or otherwise, recent approaches use diverse strategies with specific limitations. GraspGPT [6] and FoundationGrasp [7] are most similar to our method, but due to their reliance on TaskGrasp, have limited pretraining with simple scenes and tasks. LERF-TOGO [8] uses language-embedded radiance fields to localize grasp points in a scene, but has costly requirements such as multiple views and long inference times. RTAGrasp [9] transfers grasping behavior from human demonstrations but depends heavily on task-grounded data. CROG [10] and GraspCLIP [11] directly infer per-pixel predictions for grasp poses, but are limited to 4-DoF grasps, and strictly require a top-down camera view. TOGNet [12] has limited task-aware grounding, resulting in coarse semantics that are ill-suited for diverse tasks. Our work addresses these collective limitations through synthetic data generation with complex scenes, diverse and natural task specifications, and large data scale.
@@
Vision-language models in robotics Recent work has shown that leveraging foundation vision-language models (VLMs) can yield significant progress towards generalizability in robotics [13, 5, 14, 15, 16, 17, 18, 19, 20]. Within this paradigm, some directly use pre-trained VLMs [21, 22, 23, 24], while others fine-tune these models to enhance specific capabilities, including action prediction [25, 17, 19], physics [26], navigation [27, 28], and spatial reasoning [29, 30, 31]. Our method aligns with the latter approach, fine-tuning a state-of-the-art VLM [5] specifically for task-semantic grasping tasks. While other works may predict object relationships [30, 26] or general pointing [29], we specifically ground our pointing predictions in object affordances, i.e. how and where an object should be grasped. In contrast to vision-language-action models (VLAs) that directly predict low-level robot actions [20, 19, 32], our method directly outputs grasps. This avoids embodiment-specific constraints and shortcuts having to learn redundant behavior, such as free-space motion. While works like [28, 33, 29, 27] also predict points or high-level plans that aim to be generalizable, they either only focus on navigation, or don’t specifically finetune grasping affordances into the model, resulting in a lack of robustness for grasp prediction. We also consider recent works [30, 34, 28] that show that simulation is a promising and scalable data source for real-world generalization. To that end, we generate large-scale, task-oriented grasping data that focuses on complex task-dependent object manipulation, going beyond simpler QA or navigation tasks in earlier work.
@@
Object affordances in robotics. Affordances describe the functional properties of objects, i.e. how they can be manipulated, linking perception to action beyond appearance. Affordance prediction has driven advances in learning-based 6-DoF grasping [35, 36, 37] and stable object placement [38, 39, 40, 29]. Researchers have explored various representations, including part-level segmentation [41], dense visual descriptors [42], and keypoint-based encodings [43, 44, 45]. For policies, affordances are learned implicitly from demonstrations and generalized using visual features, as seen in TransporterNet [38], CLIPort [46], and recent works leveraging vision-language models (VLMs). However, many of these approaches rely on pretrained VLMs for visual grounding without incorporating deeper semantic reasoning, often resulting in less accurate or unstable point detection. In contrast, our approach fine-tunes Molmo, a VLM enhanced for semantic understanding, to generate task-aware, stable grasp points through contextualized spatial reasoning.
@@ 3.1 Generation of PRISM-Train and PRISM-Test
Our dataset employs assets from ShapeNet-Sem [47, 49] with grasps derived from ACRONYM [48]. We enumerate the 91 object classes used in data generation in Table 3. These assets are procedurally composed in a scene using SceneSynthesizer [50], and additional lighting variations and camera randomization are used to increase visual diversity. As illustrated in Figure 2, this procedure results in a comprehensive collection of scenes spanning different object categories, lighting conditions, and spatial arrangements. Furthermore, we render each scene from 10 distinct camera viewpoints to ensure that our training data captures objects from multiple perspectives. This multi-view approach is particularly important, as grasp affordances often depend on the viewpoint, which can vary wildly in the real world. We track each grasp’s visibility from each viewpoint, enabling data filtering to ensure data quality.
@@
Generating diverse task-grasp annotations at large scale is intractable, as every task-object-grasp triple must be annotated. In other words, annotating $G$ grasps per object for $O$ objects across $T$ tasks is $\mathcal{O}(TOG)$. Our insight is that we can drastically reduce the effort required to generate task-grasp annotations by exploiting common structure. Namely, we note that many different tasks might require grasping an object in the same way. Concretely, pouring from a mug and drinking from it both require grasping the handle in the same manner. Therefore, we use grasp descriptions as the bridge between tasks and grasps. Instead of individually annotating each task-object-grasp triple with a binary label, we annotate each object-grasp pair with a natural language description of how the grasp is gripping the object. Separately, for each task we generate descriptions of how an object should be grasped to complete that task. Then, matching tasks to object-grasp pairs reduces to simple text matching. The use of common-sense reasoning and natural language make this task well-suited for automation via LLMs, further aiding in scalability. This decomposition decouples tasks from object-grasps pairs, meaning generating annotations is now $\mathcal{O}(T+OG)$, which is significantly more tractable to perform at large scale.
@@
To generate high-quality grasp descriptions, we employ a two-stage approach combining LLMs with human verification. We first use GPT-4o [51] to generate synthetic grasp descriptions, given multiple viewpoints of an object being grasped. An example of this is illustrated in Fig. 2. To ensure accuracy and quality, we then engage human workers through the Prolific platform to verify the LLM-generated descriptions, and provide corrections where necessary. We found that 45% of the synthetically generated grasp descriptions were judged to be accurate while the other 55% required manual editing, illustrating the importance of human verification for high quality data. Approximately \$3,400 was spent on human data filtering.
@@
We generate a rich diversity of semantic grasping tasks in two steps, leveraging the knowledge and common-sense reasoning in existing LLMs. Given an object class, we first prompt GPT-4.1 to generate two grasp descriptions that are maximally different but still plausible for manipulation tasks, avoiding hallucinating features or descriptions which may or may not exist for a specific instance of the object class. The LLM then generates four valid semantic tasks for each grasp, while minimizing compatibility of the instruction with the alternative grasp. In total, we generate 728 unique tasks corresponding to two grasps for each of the 91 object classes. Complete prompting details are available in the appendix.
@@
Finally, we match the generated grasp descriptions for each object-grasp pair with that of the generated tasks, using GPT-4o. For each object in a scene we first determine the visible grasps and corresponding grasp descriptions. For each generated task for that object, we then take the corresponding proposed grasp description and ask GPT-4o to determine which annotated object-grasp pair, if any, describes the same grasp. If the LLM determines that one of the existing grasps in the scene is described by the generated description, the task is paired with the object-grasp to create an task-object-grasp triple, and added to the dataset.
@@
In total, PRISM-Train  contains 9,424 unique object-centric grasp instances across 2,356 unique objects, which is scaled combinatorially via procedural generation. Although we do not use this information for training GraspMolmo, the dataset also includes calibrated 3D grasp poses for all objects, providing ground truth spatial information that can be used for both 2D and 3D perception tasks. The final dataset is 378,844 samples: each has a scene render, a task in natural language, a ground-truth semantically appropriate and stable grasp in calibrated camera coordinates, an object-centric spatial description of that grasp in natural language, and a pixel location corresponding to the correct grasp point for that task. We present sample generated scenes from PRISM in Figure 4(b), showcasing the diversity in objects, arrangements, and appearances.
@@ 3.2 TaskGrasp-Image
TaskGrasp [2] is a standard dataset in the field of task-oriented grasping, used for training and evaluation. The TaskGrasp dataset consists of partial object point clouds, annotated with stable grasp poses. Each grasp on each object is annotated with binary labels, indicating whether each grasp satisfies a given task verb (e.g., scoop, cut, stir). However, the use of point clouds makes using image-based models difficult, especially due to limitations in the data such as noisy point cloud reconstructions caused by self-occlusions and segmentation artifacts. Additionally, the use of single verbs and nouns to define a task results in simplistic task conditioning, which insufficiently captures the richness of real-world human instructions.
To address these challenges, we construct TaskGrasp-Image, a new image-based dataset derived from the existing TaskGrasp dataset. Concretely, TaskGrasp-Image is made up of the original RGB-D images from TaskGrasp instead of the fused point clouds assembled from said images. This yields more realistic input due to the absence of fusion, filtering, or segmentation artifacts. As a result, TaskGrasp-Image preserves the ground-truth grasp annotations while placing them in the context of real, unprocessed RGB-D imagery. We defer complete details on deriving TaskGrasp-Image from TaskGrasp to Appendix A.6.